Curtis Products manufactures garden tools. It uses a job costing system that allocates factory overhead on the basis of direct labor-costs. Budgeted factory overhead for the year 2019 was $558,000, and management budgeted $372,000 in direct labor costs. The company’s tax rate is 21% (Round income tax to the nearest dollar).

The trial balance of Curtis Products, Inc. on June 30, 2019 is as follows:
Cash……………………………………………… $21,000
Accounts Receivable…………………………….. 65,000
Finished Goods………………………………….. 103,000
Work in Process…………………………………. 31,000
Materials ………………………………………… 19,000   
Prepaid factory Insurance.....……………………… 6,000
Machinery and Equipment (cost)……………… 264,000
Bond payable (due 05/31/2024)………………..​ $150,000
Accumulated Depreciation……………………… 70,000
Additional Paid-in Capital………………………​ 110,000
Accounts Payable……………………………….. 69,000
Common Stock……………………………………. 50,000
Retained Earnings……………………………… 60,000
Total ………………………………………… $509,000 $509,000
During July, the following transactions took place:

July 1. Purchased raw materials for $45,000 and factory supplies for $3,000 on account.
(Record materials and supplies in the materials control account)
2. Received a $50,000 loan from First National Bank to be repaid in 5 years.
3. Declared a $1,500 cash dividend.
3. Issued raw materials to production, $52,000
5. Paid factory utility bill, $10,230 in cash.
15. Received a bill for July’s janitorial services, $4,500. This amount has not yet been paid.
16.​Used factory supplies costing $1,000. (Indirect manufacturing cost).
20. Incurred research & development expense on account, $6,000.
22. Paid other factory overhead costs, $12,500.
23. Incurred selling expenses on account, $6,150 28. Paid cash dividend, $1,500.
25. Incurred payroll costs of $45,000 (not yet paid). Of this amount, $30,000 were direct labor costs
and $15,000 indirect labor cost.
27. Incurred restructuring costs totaling $15,000 on account.
27. Incurred administrative expenses on account, 7,100.
28. Prepaid factory insurance expired, $1,750.
28. Paid payroll costs, $45,000.
30. Record depreciation for July. Depreciation rate is 10% per annum on the cost of machinery and equipment. One-fifth of this depreciation is for office equipment and the remainder for factory equipment.
30. Allocate manufacturing overhead costs to production on the basis of direct labor costs.
31. Completed and transferred goods with a total cost of $95,000 to the finished goods storeroom.
31. Sold goods costing $185,000 for $374,000. (Assume all sales were made on account).
31. Paid accounts payable totaling, $158,000
31. Collected accounts receivable in the amount of $320,000.
31. Accrue $400 interest expense.
31. Calculate the overallocated or underallocated overhead and close this amount to the Cost of Goods Sold account.

Required: Prepare a multi step income statement and classified balance sheet for July 2019.
Answer #1
Spread Sheet- Accounting Equation method:
Overhead allocation rate = 558000/372000 = 1.5 per DLC
Date Assets = Liabilities + Shareholders' Equity
Cash + Acc Rec. + FG + WIP + Mat. + Fac Sup + Fac O/h + Prep F. Ins. + PPE = AP + Bond P + Loan + Acc. Dep.+ Com Sto + APIC + RE Details
OB 21000 65000 103000 31000 19000 6000 264000 = 69000 150000 70000 50000 110000 60000
1-Jul 45000 3000 = 48000
2 50000 = 50000
3 -1500 = -1500 dividend
3 52000 -52000 =
5-Fact utilities -10230 10230 =
15-janitorial 4500 = 4500
16 -1000 1000 =
20-R&D payable 6000 = 6000
22-other Fac o/h -12500 12500 =
23-selling exp. = 6150 -6150 Selling exp.
28-dividend -1500 = -1500 dividend
25-salary payable 30000 15000 = 45000
27-restructuring exp. = 15000 -15000 admin exp.
27-admin exp. = 7100 -7100 admin exp.
28 1750 -1750 =
28 -45000 = -45000
30 21120 = 26400 -5280 Admin exp.
30-O/h allocation 45000 -45000 =
31-FG 95000 -95000 =
31-Sale 374000 = 374000 Sales
31-COGS -185000 = -185000 COGS
31 -158000 = -158000
31 320000 -320000 =
31-Interest payable = 400 -400 Interest exp.
31-under allocated to COGS -27100 = -27100 COGS
TOTAL 162270 119000 13000 63000 12000 2000 0 4250 264000 = -1850 150000 50000 96400 50000 110000 184970
G TOTAL 639520 = 639520
Income Statement
Amount $
Sales 374000
Less:COGS 212100
Gross Profit 161900
Less:Admin exp. 27380
Income before selling exp. 134520
Less:Selling exp. 6150
Income before interest 128370
Less:interest exp. 400
Income before taxes 127970
Less:Taxes - 21% 26874
Net Income 101096
Add: OB of Retained Earnings 60000
Less:Dividend paid 3000
CB of Retained Earnings 158096
Balance Sheet:
Assets Amount $
Cash 162270
Acc. Receivable 119000
Inventory-FG 13000
Inventory-WIP 63000
Inventory-Material 12000
Inventory-supplies 2000
Prepaid Fac Insurance 4250
Current Assets 375520
PPE 264000
Less: Acc. Depreciation 96400
Total Assets 543120
Liabilities & Stockholders' Equity Amount $
Accounts Payable -2250
Tax payable 26874
Interest payable 400
Current Liabilities 25024
Long term liabilities:
bond payable 150000
Loans 50000
Total Liabilities 225024
Stockholders' Equity:
Common Stock 50000
APIC 110000
Retained Earnings 158096
Total Liabilities & SH. Eq. 543120